OpenStudy (anonymous):

e^x/y = 2x − y Solving by implicit

zepdrix (zepdrix):

\[\Large\rm e^{x/y}=2x-y\]

zepdrix (zepdrix):

Derivative huh? c:

zepdrix (zepdrix):

Differentiating the exponential will be a little tricky. Maybe we should start there.

zepdrix (zepdrix):

You remember your e^x derivative, yes?

OpenStudy (anonymous):

I believe so..

zepdrix (zepdrix):

lol what is it? c:

OpenStudy (anonymous):

(d/dx) e^x = e^x ? okay I dont remember :(

zepdrix (zepdrix):

Ok good good. Differentiating an exponential gives you the same thing back.

zepdrix (zepdrix):

In our problem, since the "stuff" in the exponent position is more than just an x, we have to chain rule.

zepdrix (zepdrix):

So here is how we start our problem, taking the derivative of each side with respect to x, \[\Large\rm \color{royalblue}{\left(e^{x/y}\right)'}=\color{royalblue}{\left(2x-y\right)'}\]

OpenStudy (anonymous):

so far I have, y.1-x.y/y^2=2-y ..after chain rule

zepdrix (zepdrix):

Chaining. \[\Large\rm e^{x/y}\color{royalblue}{\left(\frac{x}{y}\right)'}=\color{royalblue}{\left(2x-y\right)'}\]That's what you're getting for this chain here?

zepdrix (zepdrix):

\[\Large\rm \left(\frac{x}{y}\right)'=\frac{x'y-xy'}{y^2}=\frac{y-xy'}{y^2}\]Does the second part of your numerator have a y'? Maybe was a typo? Check that again :)

zepdrix (zepdrix):

Also with your right side,\[\Large\rm e^{x/y}\color{royalblue}{\left(\frac{x}{y}\right)'}=2-y'\]realize that anytime you take the derivative of `y with respect to x`, you're getting this y' or dy/dx popping up.

OpenStudy (anonymous):

y.1-x.y'/y^2=2-y like this ? and ohhh okay

zepdrix (zepdrix):

Any time you differentiate a `non-x` term, you'll get a prime showing up. That's how I like to think about it at least :) Ok good, don't forget your exponential though ^^

zepdrix (zepdrix):

\[\Large\rm e^{x/y}\left(\frac{y-xy'}{y^2}\right)=2-y'\]

zepdrix (zepdrix):

All of the differentiation is done at this point. (which is actually, believe it or not, the easier part of this problem). Now we have a bunch of nasty Algebra to get through in order to solve for y'.

zepdrix (zepdrix):

Where you at girl? :O You following along? Any trouble getting to that point up there?

OpenStudy (anonymous):

WAIT before you continue ! this thingy, y', we put that when y is with x, as you said before right?

OpenStudy (anonymous):

Yes Im following along, Im writing in my note book as well so I remember :)

OpenStudy (anonymous):

and where it says = 2x-y, we put y' so 2x-y',but why there as well ?

zepdrix (zepdrix):

We put that prime thing whenever we take a derivative of a `non-x` value. Example (differentiating both of these with respect to x):\[\Large\rm (x^2)'=2x\]\[\Large\rm (y^2)'\ne 2y\]We're differentiating y with respect to x, so we attach a y' to it.\[\Large\rm (y^2)'=2y~y'\]This is actually chain rule that causes this to happen. If you think about the x^2 for a sec,\[\Large\rm (x^2)'=2x~x'=2x(1)\]The derivative of x with respect to x is just 1, so we don't pay any attention to it. That step isn't necessary.

OpenStudy (anonymous):

Oh alrighty ! I think Im starting to understand

OpenStudy (anonymous):

so now we'll be solving for y' right?

zepdrix (zepdrix):

Yes.

zepdrix (zepdrix):

I would recommend multiplying both sides by y^2 as your first step. Fractions are yucky. It's nice if we can avoid them.

zepdrix (zepdrix):

\[\Large\rm e^{x/y}(y-xy')=2y^2-y^2y'\]That step ok? :o

OpenStudy (anonymous):

Yes :)

zepdrix (zepdrix):

Mmmm what next? What do you think? >.<

zepdrix (zepdrix):

sec, i gotta check on my pizza :OOO

OpenStudy (anonymous):

LOOL okay ! Ill try to figure it out while you check on your pizza :P

OpenStudy (anonymous):

Well we need to solve for y', so then hm, expnd the bracket, then solve for y'? :S

OpenStudy (anonymous):

i see two y's

OpenStudy (anonymous):

seems a bit hard

OpenStudy (anonymous):

We have to try to get the y' s on one side

zepdrix (zepdrix):

Expand the brackets, that seems like a good start.\[\Large\rm e^{x/y}y-x e^{x/y}y'=2y^2-y^2y'\]Yes, we have two y's.

zepdrix (zepdrix):

Hmm let's uhhh, add the exponential term to the other side. And we'll subtract 2y^2 from each side as well.

OpenStudy (anonymous):

got it

zepdrix (zepdrix):

Hmmmmm, what next? +_+

OpenStudy (anonymous):

take y' common, then divide the first yet by xe^x/y-y^2 ?

zepdrix (zepdrix):

Woops one of my y fell out of the problem, lemme fix that a sec D:\[\Large\rm y e^{x/y}-2y^2=x e^{x/y}y'-y^2y'\]Ok factor a y' out, then divide the bracket stuff to the other side? sounds goooood.

zepdrix (zepdrix):

\[\Large\rm y e^{x/y}-2y^2=y'(x e^{x/y}-y^2)\]

zepdrix (zepdrix):

\[\Large\rm y'=\frac{y e^{x/y}-2y^2}{x e^{x/y}-y^2}\]I'm gonna check it in Wolfram a sec, this is an easy one to make a careless mistake on.

OpenStudy (anonymous):

Alrighty !

zepdrix (zepdrix):

Oooo yay! We did it correctly! \c:/ Yayyy team!

zepdrix (zepdrix):

So on a problem like this, if you get comfortable with the y' stuff, you'll notice that the differentiation becomes fairly easy. You really need to be solid in your algebra skills to fully get through a problem like this!

OpenStudy (anonymous):

Thank you so much ! I wish you were my math prof, you explain everything wayyyy better !

zepdrix (zepdrix):

Hehe, aren't you sweet :3

zepdrix (zepdrix):

Please do give your teacher a little bit of slack though :) We spent 40 minutes on this single problem. In class you're definitely more pressed for time, so it can be difficult squeezing in every little detail that you need.

OpenStudy (anonymous):

That's true >.< But thanks again! :)